OpenStudy (anonymous):

Hong works as a tutor for $8 an hour and as a waiter for $7 an hour. This month, he worked a combined total of 100 hours at his two jobs. Let t be the number of hours Hong worked as a tutor this month. Write an expression for the combined total dollar amount he earned this month.

OpenStudy (anonymous):

8t + 7x = 100

OpenStudy (solomonzelman):

I don't think this is correct. 100 is not the total earning, it is the number of hours.

OpenStudy (solomonzelman):

\(\large\color{#CC3300}{ \displaystyle t+x }\) is the total number of hours, and thus you have: \(\large\color{#CC3300}{ \displaystyle t+x =100 }\)

OpenStudy (solomonzelman):

However this is not the total earning, rather, just the hours worked.

OpenStudy (solomonzelman):

The total earning would be: \(\large\color{#CC3300}{ \displaystyle 8t+7x }\)

OpenStudy (anonymous):

Ohh, I see. I thought it said dollars. My bad.